OpenStudy (cookiimonster627):

use half angle identities to find the exact value. given: cos theta = -2/5 and theta lies in quadrant II, find tan theta/2.

OpenStudy (anonymous):

Do you know the quad rule of what should be negative or positive?

OpenStudy (cookiimonster627):

no, im not sure

OpenStudy (anonymous):



OpenStudy (cookiimonster627):

its going to be negative?

OpenStudy (anonymous):

Quad 1: Everything is positive Quad 2: Only sin/csc are postive Quad 3: Only tan/cotan are positive Quad 4: Only cos and sec are positive

OpenStudy (cookiimonster627):

o only cot and tan would be positive

OpenStudy (anonymous):

Sin = O/H Cos = A/H Tan O/A Draw a triangle given that cos = -2/5 in quad 3

OpenStudy (anonymous):

The Quad rule can be remembered via (All students take calclus) The trig functions as: (Oscar-has a-hairy old-as-..er, arm)

OpenStudy (cookiimonster627):

oh okay

OpenStudy (anonymous):

Use cos to find tan, and make sure tan is positive. Can you do that?

OpenStudy (cookiimonster627):

i dont think so, could you do it with me step by step

OpenStudy (cookiimonster627):

@TheSmartOne could you help me. this is my last problem then im done